# Efficient and Versatile Model for Multilingual Information Retrieval of Islamic Text: Development and Deployment in Real-World Scenarios

### v1

@@ Abstract
Vera Pavlova rttl labs, UAE v@rttl.ai          Mohammed Makhlouf rttl labs, UAE mm@rttl.ai
@@ 1 Introduction
This study examines training approaches for deploying a single retrieval model across diverse MLIR settings, enabling modern search capabilities in the Islamic domain. The goal is to help users efficiently locate relevant Quranic passages in multiple languages and access the cultural and religious heritage preserved within Islamic texts, serving both scholars and the general public. We utilize the XLM-RBase model (Conneau et al., 2020), a multilingual model trained for a general domain, as a backbone for retrieval. It is known that the performance of retrieval models typically deteriorates due to domain shift (Thakur et al., 2021). As a preliminary step, we conduct a brief domain adaptation of the XLM-RBase model using a small multilingual domain-specific corpus (approximately 100M words). This short round of pre-training resulted in significant performance improvements in retrieval tasks. Moreover, to reduce the model’s size, we conduct language reduction on the XLM-RBase model, which allows us to eliminate languages that are not needed for the current deployment, resulting in more than a 50% reduction in the model’s size. We prepare eleven retrieval models using this lightweight domain-specific multilingual large language model (MLLM) by applying four different training approaches: monolingual, cross-lingual, translate-train-all, and a proposed mixed approach that combines cross-lingual and monolingual techniques. Evaluation across monolingual, cross-lingual, and multilingual retrieval scenarios demonstrates that our proposed mixed training approach produces promising results in all settings. We conduct an in-depth analysis of the potential effects of different training configurations on the embedding space and their impact on multilingual retrieval. Additionally, we discuss the advantages of deploying a single lightweight model for various potential deployment scenarios.
@@ 4.1 Datasets
For training, we use the MS MARCO (Bajaj et al., 2018), a large-scale English-language dataset widely adopted for training and evaluating dense retrieval models. It contains over 500,000 real-world queries paired with a collection of 8.8 million passages; Bonifacio et al. (2021) released machine-translated variants of MS MARCO for 13 languages, including Arabic and Russian (the collection was not translated into Urdu). Applying machine translation, we translated MS MARCO into Urdu to include this language in the experiment. For evaluation, we combined the train and development splits of the QRCD (Qur’anic Reading Comprehension Dataset) (Malhas and Elsayed, 2020) to increase the size of the test set, resulting in 169 queries used for evaluation. The answers provided are exhaustive, meaning all Qur’anic verses directly responding to the questions have been thoroughly extracted and annotated. The queries reflect contemporary, real user information needs (examples of queries are provided in the Appendix  B), so they are both valid and salient today. This pairing of complete passage coverage with today’s intents enables rigorous IR evaluation while preserving direct applicability to current user scenarios. The language of the QRCD dataset is Arabic; to evaluate in other languages, we use verified translations of this dataset to English, Russian, and Urdu. We use the Holy Quran text (Arabic), Sahih International translation (English)(Pavlova, 2023), Elmir Kuliev (Russian), and Ahmed Raza Khan (Urdu) as retrieval collections. [^2] https://tanzil.net/trans/
@@ Appendix A Appendix
Computing Infrastructure 2 x NVIDIA RTX 3090 GPU Hyperparameter Assignment number of epochs 10 batch size 8 learning rate 2e-5 weight decay 0.01
Table 9: Hyperparameters for fine-tuning on XNLI dataset.
@@ Appendix B Real User Metrics
Query ID Query 1 How is the universe created? 2 What is the purpose of life of man on earth? 3 How is the fetus formed in the womb? 4 What function of frontal lobe of the brain has? 5 How is the rain created? 6 What is the condition at the depth of the sea? 7 Why do mountains stand still on the surface of the earth? 8 Can animals communicate in their own languages ? 9 Will we be held accountable for our deeds? 10 What is Hijab? 11 Will the world come to an end and how will it happen? 12 Who was Jesus?
Table 11: Examples of real user-queries
Query ID Query 102 Who are the people of Moses (Musa)? 125 How many are the sacred months? 138 Who remained inside the belly of the whale? 152 In how many days did Allah create the universe? 216 What are the incidents related to David (Dawood), Peace Be Upon Him? 220 Who were the messengers who lived in Egypt? 227 Did Moses (Musa) live in Egypt? 257 What is the reason for the descent of Adam from Paradise? 424 Did Islam honor women? 428 Did the Quran mention the organ that the human reasons with? 421 Does the existence of different readings of the Qur’an means it had been distorted? 401 Did the Quran mention the atom?
Table 12: Examples of QRCD dataset queries

### v2

@@ Abstract
Vera Pavlova1,2, Mohammed Makhlouf1 1rttl labs  2burevestnik.ai 1v@rttl.ai, mm@rttl.ai2v@burevestnik.ai
@@ 1 Introduction
This study examines training approaches for deploying a single retrieval model across diverse MLIR settings, enabling modern search capabilities in the Islamic domain. The goal is to help users efficiently locate relevant Quranic passages in multiple languages and access the cultural and religious heritage preserved within Islamic texts, serving both scholars and the general public. We utilize the XLM-RBase model (Conneau et al., 2020), a multilingual model trained for a general domain, as a backbone for retrieval. It is known that the performance of retrieval models typically deteriorates due to domain shift (Thakur et al., 2021; Pavlova, 2023). As a preliminary step, we conduct a brief domain adaptation of the XLM-RBase model using a small multilingual domain-specific corpus (approximately 100M words). This short round of pre-training resulted in significant performance improvements in retrieval tasks. Moreover, to reduce the model’s size, we conduct language reduction on the XLM-RBase model, which allows us to eliminate languages that are not needed for the current deployment, resulting in more than a 50% reduction in the model’s size. We prepare eleven retrieval models using this lightweight domain-specific multilingual large language model (MLLM) by applying four different training approaches: monolingual, cross-lingual, translate-train-all, and a proposed mixed approach that combines cross-lingual and monolingual techniques. Evaluation across monolingual, cross-lingual, and multilingual retrieval scenarios demonstrates that our proposed mixed training approach produces promising results in all settings. We conduct an in-depth analysis of the potential effects of different training configurations on the embedding space and their impact on multilingual retrieval. Additionally, we discuss the advantages of deploying a single lightweight model for various potential deployment scenarios.
@@ 4.1 Datasets
For training, we use the MS MARCO (Bajaj et al., 2018), a large-scale English-language dataset widely adopted for training and evaluating dense retrieval models. It contains over 500,000 real-world queries paired with a collection of 8.8 million passages; Bonifacio et al. (2021) released machine-translated variants of MS MARCO for 13 languages, including Arabic and Russian (the collection was not translated into Urdu). Applying machine translation, we translated MS MARCO into Urdu to include this language in the experiment. For evaluation, we combined the train and development splits of the QRCD (Qur’anic Reading Comprehension Dataset) (Malhas and Elsayed, 2020) to increase the size of the test set, resulting in 169 queries used for evaluation. The answers provided are exhaustive, meaning all Qur’anic verses directly responding to the questions have been thoroughly extracted and annotated. The queries reflect contemporary, real user information needs, so they are both valid and salient today. This pairing of complete passage coverage with today’s intents enables rigorous IR evaluation while preserving direct applicability to current user scenarios. The language of the QRCD dataset is Arabic; to evaluate in other languages, we use verified translations of this dataset to English, Russian, and Urdu. We use the Holy Quran text (Arabic), Sahih International translation (English), Elmir Kuliev (Russian), and Ahmed Raza Khan (Urdu) as retrieval collections. [^2] https://tanzil.net/trans/
@@ 9 Conclusion
Acknowledgment
We thank the Program Chairs of the EMNLP Industry Track for their continuous support of research on underrepresented domains. We also gratefully acknowledge the Dubai AI Campus for providing the computing infrastructure that made this work possible. We extend our appreciation to our colleagues and collaborators for their support and inspiration throughout this project.
@@ Appendix B Real User Metrics
Query ID Query
1 How is the universe created?
2 What is the purpose of life of man on earth?
3 How is the fetus formed in the womb?
4 What function does the frontal lobe of the brain have?
5 How is the rain created?
6 What is the condition at the depth of the sea?
7 Why do mountains stand still on the surface of the earth?
8 Can animals communicate in their own languages?
9 Will we be held accountable for our deeds?
10 What is Hijab?
11 Will the world come to an end and how will it happen?
12 Who was Jesus?
Table 10: Examples of real user queries